OpenStudy (anonymous):

HELP!!! Graph y= -1/6 x-6

OpenStudy (anonymous):

Is the equation \[y = -x/6-6\] ?

OpenStudy (anonymous):

its y = -1/6 x - 6

OpenStudy (anonymous):

I get a straight line when I try but that isnt one of my choices

OpenStudy (anonymous):

Could it be this? http://www.wolframalpha.com/input/?i=y+%3D+-1%2F%286+x+-+6%29

OpenStudy (anonymous):

My choices

OpenStudy (danjs):

To graph a line, you just need 2 points, pick any 2 values for x and solve y

OpenStudy (danjs):

(x , y) and (x1,y1)

OpenStudy (anonymous):

The last picture is the answer, pretty sure.

OpenStudy (danjs):

Oh, they gave you choices... y = m*x+b b - is where it crosses the y axis m - the slope of the line

OpenStudy (anonymous):

Let me see if I get the same thing

OpenStudy (danjs):

yep, the line crosses y axis at -6, and for each change in -1 unit in the y, you get 6 units in the left, for slope -1/6

OpenStudy (anonymous):

I also get D!! thanks guys!!

OpenStudy (danjs):

For practice if needed, try and write the equations for the other 3 graphs

OpenStudy (danjs):

welcome